A cellphone worth P15,000 was bought 5 years ago. If it depreciates 5% per annum, how much does it cost now?


1
Expert's answer
2021-11-03T08:19:12-0400
"C(t)=P(1-r\/n)^{n(t)}"

"C(5)=15000(1-0.05\/1)^{1(5)}=11606.71"

It costs P11606.71 now.